# A Self-Improving Coding Agent

### v1

@@ 6 Safety Considerations
First, a key safety mitigation in such systems is observability: that humans can and do oversee all aspects of potentially self-improving systems. We have thus architected SICA with sophisticated yet user-friendly observability mechanisms, which provide rich insight into the chain-of-thought, actions and sub-agent invocations at each step. These include an interactive web interface for easy oversight of all events occurring in the agent system, as well as the asynchronous LLM-based overseer, which may be prompted to treat various behaviors as grounds for canceling an agent, and which may also use a different LLM to the running agent. We use these mechanisms to oversee the system at every step. Given the lack of model weight updates, which make it difficult but not impossible for the model to encode e.g. stenographic messages (Motwani et al., 2023; 2024) in the chain-of-thought, we believe this provides strong safety mitigation. That said, for recursively self-improving systems where the model weights are also updated, observability mechanisms, on their own, are clearly inadequate, and further research on safety mechanisms is clearly necessary.

### v2

@@ 6 Safety Considerations
First, a key safety mitigation in such systems is observability: that humans can and do oversee all aspects of potentially self-improving systems. We have thus architected SICA with sophisticated yet user-friendly observability mechanisms, which provide rich insight into the chain-of-thought, actions and sub-agent invocations at each step. These include an interactive web interface for easy oversight of all events occurring in the agent system, as well as the asynchronous LLM-based overseer, which may be prompted to treat various behaviors as grounds for canceling an agent, and which may also use a different LLM to the running agent. See Appendix A.2 for an example overseer judgement and the control afforded to it at each asynchronous invocation. Given the lack of model weight updates, which make it difficult but not impossible for the model to encode e.g. stenographic messages [24, 25] in the chain-of-thought, we believe this provides strong safety mitigation. That said, for recursively self-improving systems where the model weights are also updated, observability mechanisms, on their own, are clearly inadequate, and further research on safety mechanisms is clearly necessary.
@@
We do not believe there to be any significant, direct adverse societal consequences to this work. Our objectives either focus on improving the mechanics of code editing or the effectiveness of multi-step reasoning through longer-horizon coding tasks.